# Improved Zero-Shot Classification by Adapting VLMs with Text Descriptions

### v1

@@ 5.5 Evaluation on Novel Tasks
We evaluate our model on tasks other than category identification. We use the NeWT [35] benchmark for this. We evaluate on tasks including identifying age, attribute, health, photo quality, species, context and behavior. We evaluate the model and baseline CLIP using Mean Average Precision and average over all tasks. Our model trained on iNaturalist using visual + taxonomy + location information outperforms baseline CLIP: 60.25 vs 61.90 MAP $\rightarrow$ 4.1% relative error reduction. We present all tasks and texts we use to evaluate as well as MAP per task in the Appendix. Below is an example prediction of our model.
@@ 5.6 Resource Requirements
We fine-tune our model using 1024 batch size on a single NVIDIA A100 80GB. We need to train only for a maximum of 15 epochs. For the smaller datasets such as CUB, FGVC Aircrafts, Stanford Cars this takes less than 5 minutes. For iNaturalist which is a large dataset we train only for 5 epochs, which takes about 4 hours. Even though we are fine-tuning the encoders of CLIP, we are able to achieve significantly better performance in reasonable time since we only need to train for a few epochs for the model to learn image-attribute associations.
The cost of using the GPT4-0613 API to query text descriptions for a dataset with 100-200 classes (such as CUB200, FGVC Aircrafts, Stanford Cars) is about \$1-\$5. This is low because it scales with number of categories and not the number of images. Generating captions per image is both time consuming and expensive, an estimate for doing this for the CUB dataset using the GPT4 Vision API is more than \$100. This cost scales significantly for larger datasets with more images.
@@ 5.7 Alternate Training Strategies
Here we discuss the multiple other training strategies we try, that can intuitively be used to enhance category-level VLM training. However, we find that using our presented simple fine-tuning strategy by stochastically pairing images with texts within categories, is scalable, efficient and offers similar improvement compared to more complex training strategies.
Firstly, since we pair a given image of a category to every text of that category, we might be pairing texts that describe a part of the category that is not visible or is obscured in said image. For CUB dataset we have the ground truth visibility annotations of various bird parts, which we use to neglect texts which are occluded for each image. This strategy offers improvement to our scores (54.23 $\rightarrow$ 54.47). However, this depends on ground truth visibility annotations or another model to predict visibility and thus is not scalable and time-consuming.
@@
We test the recently released GPT4 Vision API for checking quality of image captions obtained. Even though it performs better than previous captioning models such as LLaVA [15] and BLIP2 [14], we find that the captions obtained are general descriptions without fine-grained details. The captions are specific to the image but do not describe information helpful to identify the category. An example is
@@
for this image of a Vesper Sparrow, where GPT4 provides a general description of the bird and suggests the presence of clear blue sky which is not visible in the image. We provide more detail including prompt and examples in Appendix.
@@ 6 Limitation
Since our method is trained on texts generated by LLMs, it is important to verify the correctness of these texts to assess the level of noise in our training dataset. As described in § 3.1, we conduct spot checks across many categories of CUB and our evaluation sets consists of popular categories where we expect high accuracy from LLMs. For larger datasets, however, it becomes impractical to check enough categories to fully gauge the LLM’s domain knowledge. Our method is predicated on the ability to learn meaningful information, even when attributes within the training data are somewhat noisy. Evaluation on human generated texts on CUB supports this.
@@ 7 Conclusion
We present a method to improve the zero-shot performance of VLMs using attributes generated by LLMs on fine-grained domains. Our evaluation strategy involves testing the trained model on unseen classes, texts generated from different LLMs as well as humans, dissimilar domains, and novel tasks. We show that fine-tuning CLIP using category-level descriptions from GPT4 significantly improves performance compared to baselines in this challenging downstream evaluation framework. Our findings suggests that habitat and geographic priors are equally effective and complementary to visual information for zero-shot classification in natural domains. We plan to publicly release our benchmark across all 7 datasets.
@@ 8 Acknowledgements
We thank Aaron Sun, Gustavo Perez, Max Hamilton, Rangel Daroya, and Mustafa Chasmai for participating in the verification of attributes generated by GPT4. The project is supported in part by NSF Grants #2329927 and #1749833. Our experiments were performed on the GPU cluster funded by the Mass. Technology Collaborative.
@@ Appendix A Category-level Texts produced by LLMs
We provide more examples of texts produced for different prompts, including both visual and habitat. We also list the ‘System Role’ which is high-level instruction of behavior of the LLM. For prompting about visual attributes we set ‘System Role’ as “You are a helpful assistant who can identify any [domain] based on its photo.” For prompting about habitat attributes we set ‘System Role’ as “You are a helpful assistant who knows about the habitat of any [domain].”
@@ A.1 GPT4 texts used in our method
a type of plant with scientific name Picea glauca based on just a photo?
@@
Texts should be of the form “Plant with <description>”.
@@
Each text produced should start with “Plant which”.
@@
Figure 4: Example of texts produced for a category “White Spruce” of the iNaturalist dataset using GPT4-0613, example images on the right.
@@
Texts should be of the form “a bird with <characteristic feature>”.
@@
Each text produced should start with “a bird”.
@@
Figure 5: Example of texts produced for a category “Common Tern” of the CUB dataset using GPT4-0613, example images on the right.
@@ A.2 LLaMA texts used during evaluation
Texts should be of the form “a bird with <characteristic feature>”.
@@
Each text produced should start with “a bird”.
@@
Table 6: Example of texts produced for category “Common Tern” of the CUB dataset using LLaMA2-7B.
@@ Appendix C Training Hyperparameters
CUB Stanford Cars Flowers102 FGVC Aircraft NABirds iNat EuroSAT Food101 ImageNet
lr proj 6E-07 7E-07 7E-07 3E-07 8E-07 1E-07 2E-06 2E-07 2E-06
lr main 1E-07 1E-07 1E-07 1E-07 2E-07 5E-08 7E-07 5E-08 7E-07
wd proj 1E-01 1E-02 1E-02 1E-06 1E-06 1E-02 1E-06 1E-06 1E-06
wd main 1E-02 1E-03 1E-03 1E-06 1E-06 1E-03 1E-06 1E-06 1E-06
temperature init 1.3 1.6 2.0 1.8 1.0 1.0 1.8 2.0 1.8
@@ Appendix F Improvement over CLIP-A-Self [17]
CLIP-A-self only tunes a adapter network and lags in performance compared to ours (CLIPFT + A). This again points to the need for fine-tuning CLIP encoders to recognize fine-grained attributes. Another reason why CLIP-A-self suffers is because they query GPT to produce descriptive texts for a fixed set of pre-defined attributes. It should also be noted that CLIP-A-self performs worse than CLIP-G in most cases, which shows that even though it does not tune the CLIP encoders, it still overfits to the training classes.
@@ Appendix G Correctness of LLM generated texts
Table 8: Correctness study of the category “Downy Woodpecker” of the CalTech-UCSD Birds dataset
@@
Table 9: Correctness study of the category “Spitfire” of the FGVC Aircraft dataset
@@
Table 10: Correctness study of the category “Rolls-Royce Ghost Sedan 2012” of the Stanford Cars dataset

### v2

@@ 5.5 Evaluation on Novel Tasks
We use the NeWT [40] benchmark to evaluate on tasks beyond categorization. These include identifying age, attribute, health, photo quality, species, context and behavior. We evaluate the model and baseline CLIP using average Mean Average Precision (MAP) across tasks. Our model trained on iNaturalist using visual + taxonomy + location information outperforms baseline CLIP: 60.25 vs 61.90 MAP $\rightarrow$ 4.1% relative error reduction. We present all tasks and texts we use to evaluate as well as MAP per task in the Appendix. Below is an example prediction:
@@ 5.6 Resource Requirements
We fine-tune our model using a 1024 batch size on a single NVIDIA A100 80GB. We need to train only for a maximum of 15 epochs. For smaller datasets such as CUB, FGVC Aircraft, and Stanford Cars, this takes less than 5 minutes. For iNaturalist, which is a large dataset, we train only for 5 epochs, taking about 4 hours.
The cost of using the GPT-4 API to query text descriptions for a dataset with 100-200 classes (such as CUB) is about \$1-\$5. This is low because it scales with the number of categories and not the number of images. Generating captions per image is both time-consuming and expensive; an estimate for doing this for the CUB dataset using the GPT-4 Vision API is more than \$100. This cost scales significantly for larger datasets with more images.
@@ 5.7 Alternate Training Strategies
Here we discuss other training strategies for improving VLMs using category-level training data. Our simple fine-tuning strategy of stochastically pairing images with texts within categories is simple, efficient, and offers similar improvements compared to more complex approaches.
Firstly, since we pair a given image of a category to every text of that category, we might be pairing texts that describe attributes that are not visible. For CUB dataset we have the ground truth visibility annotations of various bird parts, which we use to ignore texts that are occluded for each image. This strategy offers improvement to our scores (54.23 $\rightarrow$ 54.47). However, this depends on visibility information that is time-consuming to generate.
@@
We test the recently released GPT4 Vision API for checking quality of image captions obtained. Even though it performs better than previous captioning models such as LLaVA [17] and BLIP2 [16], we find that the captions obtained are general descriptions without fine-grained details. The captions are specific to the image but do not describe information helpful to identify the category. An example is:
@@
For this image of a Vesper Sparrow, GPT4 provides a general description of the bird and suggests the presence of clear blue sky which is not visible in the image. We provide more detail including prompt and examples in Appendix.
@@ 6 Limitation
Since our method is trained on texts generated by LLMs, it is important to verify the correctness of these texts to assess the level of noise in our training dataset. As described in § 3.1, we conduct spot checks across some categories on our evaluation sets. However, for our larger training datasets, vetting becomes impractical. Improved performance on human-generated texts on CUB, as well as the vetted evaluation sets, supports our model’s ability to learn meaningful information from somewhat noisy training data.
@@ 7 Conclusion
We present a method to improve the zero-shot performance of VLMs using attributes generated by LLMs on fine-grained domains. Our evaluation strategy involves testing the trained model on unseen classes, texts generated from different LLMs as well as humans, dissimilar domains, and novel tasks. We show that fine-tuning CLIP using category-level descriptions from GPT4 significantly improves performance compared to baselines in this challenging downstream evaluation framework. Our findings suggests that habitat and geographic priors are equally effective and complementary to visual information for zero-shot classification in natural domains. We publicly release our benchmark across all 14 datasets.
@@ 8 Acknowledgements
We thank Aaron Sun, Gustavo Perez, Rangel Daroya, and Mustafa Chasmai for participating in the verification of attributes generated by GPT4. The project is supported in part by NSF Grants #2329927 and #1749833. Our experiments were performed on the GPU cluster funded by the Mass. Technology Collaborative.
@@ Appendix A Category-level Texts produced by LLMs
We provide more examples of texts produced for different prompts, including both visual and habitat. For prompting about visual attributes we set ‘System Role’ as “You are a helpful assistant who can identify any [domain] based on its photo." For prompting about habitat attributes we set ‘System Role’ as “You are a helpful assistant who knows about the habitat of any [domain]."
@@ A.1 GPT4 texts used in our method
a type of plant with scientific name Picea glauca from other organisms based on just a photo?
@@
Texts should be of the form “Plant with <description>".
@@
Each text produced should start with “Plant which".
@@
Figure 4: Example of texts produced for a category “White Spruce" of the iNaturalist dataset using GPT4-0613, example images on the right.
@@
Texts should be of the form “a bird with <characteristic feature>".
@@
Each text produced should start with “a bird".
@@
Figure 5: Example of texts produced for a category “Common Tern" of the CUB dataset using GPT4-0613, example images on the right.
@@ A.2 LLaMA texts used during evaluation
Texts should be of the form “a bird with <characteristic feature>".
@@
Each text produced should start with “a bird".
@@
Table 6: Example of texts produced for category “Common Tern" of the CUB dataset using LLaMA2-7B.
@@ Appendix C Training Hyperparameters
CUB FGVC Aircraft Stanford Cars Flowers102 NABirds iNat EuroSAT Food101 ImageNet CalTech101 DTD Oxford Pets Sun397 UCF101
lr proj 6E-07 4E-07 3E-07 7E-07 8E-07 1E-07 2E-06 1E-07 2E-06 1E-07 5E-07 8E-07 1E-06 8E-07
lr main 1E-07 1E-07 1E-07 1E-07 1E-07 5E-08 5E-07 5E-08 5E-07 5E-08 2E-07 2E-07 5E-07 2E-07
wd proj 1E-01 1E-06 1E-02 1E-02 1E-06 1E-03 1E-06 1E-06 1E-06 1E-03 1E-06 1E-04 1E-06 1E-06
wd main 1E-01 1E-06 1E-02 1E-03 1E-06 1E-03 1E-06 1E-06 1E-06 1E-03 1E-06 1E-04 1E-06 1E-06
temperature init 1.3 1.8 1.6 2 1 1 1.8 2 1.8 2 2 2 2 2
@@ Appendix F Improvement over CLIP-A-Self [19]
CLIP-A-self only tunes a adapter network and lags in performance compared to ours (CLIPFT + A). This again points to the need for fine-tuning CLIP encoders to recognize fine-grained attributes. Another reason why CLIP-A-self suffers is because they query GPT to produce descriptive texts for a fixed set of pre-defined attributes. It should also be noted that CLIP-A-self performs worse than CLIP + A in most cases, which shows that even though it does not tune the CLIP encoders, it still overfits to the training classes. We provide results on the all 14 datasets in Tab. 8. For B/16 we compare with CLIP-A-Self in a 16-shot setting for 1:1 train/test split. In Tab. 9 we also compare with 3:1 train/test split for CUB dataset.
Training Dataset CUB FGVC Aircraft Stanford Cars Flowers102 NABirds iNat EuroSAT Food101 ImageNet CalTech101 DTD Oxford Pets Sun397 UCF101
Testing Dataset CUB CUB Flowers
B/32 CLIP 50.54 29.27 69.72 71.78 50.54 50.54 71.78 68.89 89.94 62.61 93.87 58.45 96.64 73.03 71.12
B/32 CLIP + A 50.71 30.35 69.47 75.37 50.71 50.71 75.37 72.07 90.06 65.11 94.54 59.90 96.53 75.97 75.71
B/32 CLIPFT + A 53.34 36.41 71.63 77.05 55.29 54.58 77.05 78.56 93.71 65.98 95.75 62.20 96.88 77.75 75.99
B/16 CLIP 51.91 36.47 74.94 77.05 51.91 51.91 77.05 64.05 92.49 67.41 93.89 60.26 97.04 75.49 77.45
B/16 CLIP + A 53.58 36.47 73.83 80.84 53.58 53.58 80.84 71.51 93.72 69.74 94.87 64.13 96.81 78.88 80.47
B/16 CLIP-A-Self - 33.00 72.90 75.30 - - 75.30 70.50 91.20 68.30 95.90 62.30 97.00 76.80 76.40
B/16 CLIPFT + A 55.63 40.75 75.78 81.26 56.76 56.77 81.20 81.82 95.08 71.87 96.03 65.21 97.21 80.32 80.69
Table 8: Results on 14 datasets. We show performance on ViT B/32 and B/16 with 1:1 train/test class split, obtaining superior accuracy across all settings.
Method Accuracy
CLIP 71.18
CLIP + A 72.54
CLIPFT + A 74.80
CLIP-A-Self 71.30
Table 9: Performance on CUB with 3:1 train/test split. Following CLIP-A-Self CUB split we show that we outperform significantly.
CUB FGVC Aircraft Stanford Cars
Multiply Probs 53.10 36.35 71.58
Average Probs 53.34 36.41 71.63
Table 10: Averaging vs Multiplying Probabilities for aggregating. Averaging works better than multiplying per-text probabilities for each class for final classification accuracy.
@@ Appendix G Correctness of LLM generated texts
Table 11: Correctness study of the category “Downy Woodpecker" of the CalTech-UCSD Birds dataset
@@
Table 12: Correctness study of the category “Spitfire" of the FGVC Aircraft dataset
@@
Table 13: Correctness study of the category “Rolls-Royce Ghost Sedan 2012" of the Stanford Cars dataset